Find three consecutive odd integers such that the sum of the last two is 15 less than five times the first
Answer by mananth(16946)

Find three consecutive odd integers such that the sum of the last two is 15 less than five times the first
let the integers be x , x+2, x+4
the sum of the last two= x+2+x+4 = 2x+6
5 times first = 5x
15 less than 5times first = 5x-15
equate
2x+6 = 5x-15
solve
3x=21
/3
x=7
the numbers are 7,9,11